# Adaptive Non-uniform Timestep Sampling for Accelerating Diffusion Model Training

### v1

@@ 6 Experiments
Frechet Inception Distance (FID) [13] is used to evaluate both the fidelity and coverage of generated images. We mainly used a machine equipped with four NVIDIA RTX 3090 GPUs to train the models. Detailed hyperparameter values and implementation details can be found in Appendix Sec. 9 and Sec. 10.
@@ 6.1 Comparison to Heuristic Acceleration Methods
To compare our model and other acceleration techniques across various noise schedules, we used the CIFAR-10 32$\times$32 dataset [2], utilizing DDPM [14] as the backbone architecture.
@@
To evaluate the acceleration methods on high-resolution images, we used the CelebA-HQ 256$\times$256 dataset [16], utilizing LDM [24] as the backbone architecture. Tab. 2 demonstrates that although heuristic methods initially achieve better performance, our learning-based approach surpasses them over time. We also found that P2 performs comparably well in this setting relative to the results in Tab. 1, suggesting that the performance of heuristics can vary across different settings. Visualizations of the generated CelebA-HQ images from our method and baselines are in Appendix Sec. 11.
@@ 6.2 Compatibility with Heuristic Approaches
To assess the compatibility of our model with heuristic approaches, we trained our model along with Min-SNR weighting on the ImageNet 256$\times$256 benchmark [8]. We used the encoder from LDM [24] to convert images of dimensions $256\times 256\times 3$ into latent representations of size $32\times 32\times 4$. We used U-ViT [5] as our diffusion backbone. Tab. 3 shows that our method with heuristic weighting can achieve faster convergence from the beginning as well, ultimately converging to a lower FID score compared to the baseline.
@@ 9 Implementation Details
Dataset CIFAR-10 32$\times$32 CelebA-HQ 256$\times$256 ImageNet 256$\times$256
@@
Dataset CIFAR-10 32$\times$32 CelebA-HQ 256$\times$256 ImageNet 256$\times$256
@@ 10 Hyperparameter settings
Timestep Sampler Learning Rate 1e-2
@@
Timestep Sampler Optimizer Adam $(\beta_{1}=0.9,\beta_{2}=0.999)$
Learning Rate 1e-3
@@
Timestep Sampler Optimizer AdamW $(\beta_{1}=0.99,\beta_{2}=0.99)$
Learning Rate 1e-3

### v2

@@ 6 Experiments
Frechet Inception Distance (FID) [13] is used to evaluate both the fidelity and coverage of generated images. We mainly used a machine equipped with NVIDIA RTX 4090 GPUs to train the models. Detailed hyperparameter values and implementation details can be found in Appendix Sec. 10 and Sec. 11.
@@ 6.1 Comparison to Heuristic Acceleration Methods
To compare our model and other acceleration techniques across various noise schedules, we used the CIFAR-10 $\times32$ dataset [2], utilizing DDPM [14] as the backbone architecture.
@@
To evaluate the acceleration methods on high-resolution images, we used the CelebA-HQ $\times256$ dataset [16], utilizing LDM [24] as the backbone architecture. Tab. 2 demonstrates that although heuristic methods initially achieve better performance, our learning-based approach surpasses them over time. We also found that P2 performs comparably well in this setting relative to the results in Tab. 1, suggesting that the performance of heuristics can vary across different settings. Visualizations of the generated CelebA-HQ images from our method and baselines are in Appendix Sec. 12.
@@ 6.2 Compatibility with Heuristic Approaches
To assess the compatibility of our model with heuristic approaches, we trained our model along with Min-SNR weighting on the ImageNet $\times256$ benchmark [8]. We used the encoder from LDM [24] to convert images of dimensions $256\times 256\times 3$ into latent representations of size $32\times 32\times 4$. We used U-ViT [5] as our diffusion backbone. Tab. 3 shows that our method with heuristic weighting can achieve faster convergence from the beginning as well, ultimately converging to a lower FID score compared to the baseline.
@@ 7 Conclusion
8 Acknowledgements
This work was partly supported by Institute of Information &\& communications Technology Planning &\& Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II220311, Development of Goal-Oriented Reinforcement Learning Techniques for Contact-Rich Robotic Manipulation of Everyday Objects, No. RS-2024-00457882, AI Research Hub Project, and No. RS-2019-II190079, Artificial Intelligence Graduate School Program (Korea University)), the IITP(Institute of Information &\& Coummunications Technology Planning &\& Evaluation)-ITRC(Information Technology Research Center)(IITP-2024-RS-2024-00436857) grant funded by the Korea government (Ministry of Science and ICT), the NRF (RS-2024-00451162) funded by the Ministry of Science and ICT, Korea, BK21 Four project of the National Research Foundation of Korea, and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(RS-2025-00560367).
@@ 10 Implementation Details
Dataset CIFAR-10 $\times32$ CelebA-HQ $\times256$ ImageNet $\times256$
@@
Dataset CIFAR-10 $\times32$ CelebA-HQ $\times256$ ImageNet $\times256$
@@
Feature Selection method
To overcome the computational burden of calculating $\delta_{k,i}^{t}$ across all timesteps, we employ a feature selection method. Specifically, we treat $\delta^{t}_{k,i}$ as features to predict the target $\Delta_{k}^{t}$ using a linear regression model. The F-statistic is then computed for each feature $\delta^{t}_{k,i}$ and by summing the top $M$ features with the highest F-statistics, we can efficiently approximate $\Delta_{k}^{t}$ while focusing on the most influential timesteps $i$. The F-statistic identifies the features that have the strongest linear relationship with the target variable, enabling us to focus on the most critical timesteps in the approximation process. Although the F-statistic is employed here, other feature selection methods could also be applied.
@@ 11 Hyperparameter settings
Timestep Sampler Learning Rate 1e-2(linear, quad), 1e-3(cosine)
@@
Timestep Sampler Learning Rate 1e-3
@@
Timestep Sampler Learning Rate 1e-3